use partial derivative to solve for uf(x,y,t)

if f(x,y,t) is 5X^3y-1/xy+1/xt+x^2t^2

Ux=0.5

Uy=0.2

Uz=0.3

Use numerical methos to solve also this problem
Answer #1

ANSWER 8 Usinq poatial derivateve - szý- way to 2 V txt xt film,4,+) = 57y- to solve Of(x,y,t) = ? Ux=0.5 0g •0-2 Uz = 0-3 Uf